# VividDreamer: Invariant Score Distillation For Hyper-Realistic Text-to-3D Generation

### v1

@@ 0.A.1 Implementation Details
0.A.2 Visualization of Over-saturation
Figure 9: In order to show the relationship between over-saturation and guidance scale $w$ in SDS, we use different guidance scales for text-to-image generation. We find that when $w$ increases, the over-saturation problem of the generated image gradually becomes serious. Therefore, to avoid the generated samples suffering from over-saturation, we should use a smaller guidance scale.
In Fig. 2, we show that using only the classifier-free guidance term leads to the over-saturation problem. In this section, we further demonstrate the quality of generated images under different guidance scales using SDS [37] to reveal the relationship between over-saturation and classifier-free guidance scale $w$.
As can be seen in Fig. 9, we find that as $w$ increases, the images generated by SDS gradually become full of details (such as the texture of the horse, the details of the astronaut), but the over-saturation problem is also gradually amplified. Therefore, we speculate that the over-saturation phenomenon in SDS comes from the larger guidance scale. This phenomenon is also reasonable because Stable Diffusion [40] has not used classifier-free guidance term alone for generating during pre-training. Therefore, this item itself cannot generate real samples, and amplifying its weight will have a negative impact.
0.A.3 Visualization of Over-smoothing
@@
0.A.4 SDS fail in a normal guidance scale
@@
0.A.5 Quantify the Over-saturation
To compare the saturation among methods, we calculated the saturation and brightness of 50 generated samples(Tab. LABEL:table:more_eva). The average saturation of ISD is 29% (203.75 to 144.92) lower than that of SDS. In contrast, the average brightness of ISD is higher than that of SDS.
@@ Appendix 0.D Additional Results
Furthermore, in Fig. LABEL:fig:more_comp,fig:more_comp2,fig:more_comp3, we show additional comparisons of the results generated by our method with other methods.

### v2

@@ 6 Conclusion
Acknowledgments
This work is supported by the Major program of the National Natural Science Foundation of China (T2293720/T2293723), the National Natural Science Foundation of China (U2336212), the National Natural Science Foundation of China (62293554), the Fundamental Research Funds for the Central Universities (No. 226-2024-00058), the Fundamental Research Funds for the Zhejiang Provincial Universities (226-2024-00208), and the China Postdoctoral Science Foundation (524000-X92302).
@@ 0.A.1 Implementation Details
0.A.2 Visualization of Over-smoothing
@@
0.A.3 SDS fail in a normal guidance scale
@@
Table 3: More evaluation
Methods Average Average
saturation brightness
SDS 203.75 171.74
NFSD 170.60 190.14
VSD 153.51 207.91
ISD 144.92 216.97
Table 4: More ablation on $c$
[…]
Score ()($\uparrow$)( ↑ ) R@1 R@5 R@10
ISD (c=0) 73.37 75.69 90.07 96.53
ISD (c=1) 76.59 79.93 93.35 97.42
ISD (c=20) 81.06 82.30 95.08 97.93
ISD (c=50) 80.87 82.44 94.98 97.79
0.A.4 Quantify the Over-saturation
To compare the saturation among methods, we calculated the saturation and brightness of 50 generated samples(Tab. 4). The average saturation of ISD is 29% (203.75 to 144.92) lower than that of SDS. In contrast, the average brightness of ISD is higher than that of SDS. This further proves the effectiveness and superiority of the ISD.
0.A.5 Quantitative ablation study on hyperparameter $c$
The hyperparameter $c$ is related to the richness of detail in the generated results. We provide more quantitative results in Tab. 4. We find that a smaller $c$ results in a lower CLIP Score due to the loss of details, whereas a conventional value of $c$ yields better results.
@@ Appendix 0.D Additional Results
Furthermore, in Fig. 14, 15, 16, we show additional comparisons of the results generated by our method with other methods.